OpenStudy (anonymous):

A point on the graph of y = 1/x is moving along the curve in such a way that its x-coordinate is increasing at a rate of 3 units per second. What is happening to the y-coordinate at the instant the y-coordinate is equal to 2

OpenStudy (anonymous):

when y=2, 2=1/x x=1/2 y=1/x \[\frac{ dy }{dt }=\frac{ -1 }{ x ^{2} }\frac{ dx }{dt },\frac{ dx }{dt }=3,x=\frac{ 1 }{2 }\] \[Find \frac{ dy }{ dt }\]

OpenStudy (anonymous):

Thank you !!

OpenStudy (anonymous):

yw